SUMMER 2008 UNIVERSITY OF CHICAGO REU PROGRAM 8 week program

CHECK ONE:

I will participate in the program; I understand that it is my responsibility to be present M-F 9:00-5:00 from June 23 to August 15, 2008 and that I will not engage in any other employment during that period without the explicit approval of the program director. I also understand that a final piece of mathematical writing (expository or research) is required.

I will not participate in the program

NAME:

ADDRESS:

E-MAIL: